OpenStudy (anonymous):

Charles has been asked to insert a trapezoid in the floor tiling of the rotunda of the courthouse. One base must be 8 feet and the height must be 10 feet. Write an equation for the area.

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

algebra 1 equations

OpenStudy (anonymous):

http://www.mathopenref.com/trapezoidarea.html

OpenStudy (anonymous):

so, area = (8 feet + x feet / 2) * 10

OpenStudy (anonymous):

\[a=(\frac{ 8+x }{ 2 }) * 10\]

OpenStudy (anonymous):

If Charles wants the area of the trapezoid to be 65 square feet, what should be the measure of the second base?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

just plug 65 into a, sorry I was AFK

OpenStudy (anonymous):

65=

OpenStudy (anonymous):

\[65 = ((8+x)/2)∗10\]

OpenStudy (anonymous):

x will be the second side

OpenStudy (anonymous):

so idk what to do first

OpenStudy (anonymous):

so whats the first step

OpenStudy (anonymous):

Solve for x: 65 = 5 (8+x) Divide both sides by a constant to simplify the equation. Divide both sides of 65 = 5 (8+x) by 5: 8+x = 65/5 Use long division to simplify 65/5. 5 | 1 | 3 | 6 | 5 - | 5 | | 1 | 5 - | 1 | 5 | | 0: 8+x = 13 Isolate terms with x to the left hand side. Subtract 8 from both sides: x+(8-8) = 13-8 Look for two terms that sum to zero. 8-8 = 0: x = 13-8 Evaluate 13-8. 13-8 = 5: Answer: x = 5